Length of the rectangle exceeds its breadth by 7cm. if the length is decreased by 4cm. the area of the new rectangle is same as area of the original rectangle. find the length of original rectangle.
Answer by ankor@dixie-net.com(22740)

Length of the rectangle exceeds its breadth by 7cm.
if the length is decreased by 4cm.
the area of the new rectangle is same as area of the original rectangle.
find the length of original rectangle.
:
This does not seem possible, but assume it is
:
Let x = the width
then
(x+7) = the length
and
x(x+7) = the area
:
"if the length is decreased by 4cm. the area of the new rectangle is same as area of the original rectangle."
x(x+3) = the new area
therefore
x(x+7) = x(x+3)
divide both sides by x
x + 7 = x + 3; no sense here